OpenStudy (anonymous):

Sally bought x pounds of apples and y pounds of pears. Three times the number of pounds of pears she bought was 3 more than the number of pounds of apples she bought. The total number of pounds of apples and pears she bought was 5 pounds. Which graph best shows the number of pounds of apples and pears Sally bought? Im not very good at finding the equation its asking me, but after i get that im good

OpenStudy (hoblos):

Three times the number of pounds of pears=3 more than the number of pounds of apples 3y = 3 + x The total number of pounds of apples and pears she bought = 5 pounds x+y = 5 so the equations are 3y = 3+x x+y = 5

OpenStudy (anonymous):

thank you! :D could you help me with one more?

OpenStudy (hoblos):

yes of course i will try to

OpenStudy (anonymous):

Bess is trying to solve a puzzle where she has to figure out two numbers, x and y. The sum of twice x and 3 times y is less than 6. Also, y must be greater than 1 less than twice the x. I think one of the equations is 2x + 3y < 6 im not sure

OpenStudy (hoblos):

sum : + twice x : 2x 3 times y : 3y is less: < thus 2x + 3y < 6 greater : > 1 less than twice x : 2x-1 thus y > 2x -1

OpenStudy (anonymous):

oh i get it, thank you! you really helped me out!

OpenStudy (hoblos):

glad to help :D